# Training Consistency Models with Variational Noise Coupling

### v1

@@ 7.1 Baselines and models
As baselines, we re-implement the iCT and ECM models, corresponding to our iCT-VE and ECM-VE. As an additional model, we add CT with the minibatch Optimal Transport Coupling (-OT) proposed in (Pooladian et al., 2023; Tong et al., 2023) and used in (Issenhuth et al., 2024), to compare the effectiveness and scalability of the OT coupling with the learned one. Finally, we combine the baselines with our proposed Variational Coupling (-VC). For the models with learned coupling, we use gradient clipping with a large value (200 in all the experiments) to avoid instabilities at the early stages of training.
@@ 7.2 Ablation for different $\beta$
To see the effect of $\beta$ on the generation performance, we compare the results for different values on the FashionMNIST dataset for iCT-VE-VC in Table 1. As expected, with small values of $\beta$, the coupling distribution deviates from the sampling distribution and the performance degenerates, while increasing $\beta$ to high values reduces the benefits of the learned coupling.
@@
Table 1: Comparison of FID performance (lower is better) for one and two sampling steps, for varying values of $\beta$. The models are iCT-VE-VC and trained on the FashionMNIST dataset with the same settings described in appendix C.2. Best entries in bold.
@@
ECM-VE-VC (ours) 3.26 / 2.02 5.47 / 4.16 5.08 / 3.15
ECM-LI-VC (ours) 3.39 / 2.09 5.57 / 4.29 4.93 / 3.07
Table 3: Comparison of FID (lower is better, reported as 1-step / 2-step performance) for different models based on ECM. The model marked with a * is the baseline as reported in (Geng et al., 2024). All the other models are from our re-implementation. The best entries are highlighted in bold. For ImageNet, the results marked with $\dagger$ are obtained with models trained for $100k$ iterations, while the others use $200k$ iterations.
@@ 7.3 Results
CIFAR-10: For all the CIFAR-10 experiments, we use the DDPM++ architecture from (Song et al., 2021b) as implemented in (Karras et al., 2022), with EMA rate $0.9999$ as in (Geng et al., 2024). While this differs from the settings in (Song & Dhariwal, 2024), we found it to work better in our re-implementation. The remaining hyperparameters are the same as used in the respective baselines. From the results, we can see how using the learned coupling results in improved performance for both one and two steps generation, outperforming all the re-implemented baselines. The 1-step result from the original iCT is superior to our model. However, to the best of our knowledge, there is no open-source implementation that can reproduce the results reported in the paper. The learned coupling outperforms the minibatch OT coupling in all cases, as it is less affected by the effective (per device) batch size and the data dimensionality. Finally, our 2-step sampling performance for ECM-VE-VC is on par with the current SoTA achieved by other methods with similar settings (Wang et al., 2024; Lee et al., 2024b; Lu & Song, 2024). We empirically compare the variance of the gradients for iCT-VE and iCT-VE-VC, and show in Figure 3 how the resulting reduced variance corresponds to improved FID score. In particular, in early training the model with VC exhibits higher variance, which can be attributed to the fact that the encoder is still learning the coupling. As training continues, the coupling becomes effective at providing better data-noise pairs to the model, which results in reduced gradient variance and improved generative performance.
@@
ImageNet $64\times 64$ (class conditional): As a baseline, we reuse the settings from ECM with the EDM2-S architecture and batch size $128$. While the baseline is trained for $100k$ iteration, we found that our models with Variational Coupling needed more time to converge properly, as the encoder weights are not pretrained and initialized at random. We therefore train our re-implemented baselines and models for $200k$ iterations instead. In this case, the models with Variational Coupling outperform the other models, with the LI kernel obtaining the best overall FID, while the OT coupling performs poorly due to the small batch size and high data dimensionality. In Figure 2 we compare samples from ECM-LI and ECM-LI-VC, where we can see how the images generated with VC are more clear and detailed.
@@ Appendix B Consistency Lower Bound
In this section we prove the result from equation 8, where we show how CT with variational noise coupling is also a lower bound of the evidence. To begin with, we start by rewriting the evidence lower bound derivation from VAEs:
@@
$\leq\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{x}_{0})}\left[\log\frac{p_{\bm{\theta}}(\bm{x}_{0},\bm{z})}{q_{\bm{\phi}}(\bm{z}|\bm{x}_{0})}\right]$ (43)
$\mathcal{L}(\bm{x}_{0};\bm{\theta},\bm{\phi})$ $=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{x}_{0})}\left[\log p_{\bm{\theta}}(\bm{x}_{0}|\bm{z})\right]-\text{KL}(q_{\bm{\phi}}(\bm{z}|\bm{x}_{0})\|p(\bm{z}))$ (44)
where $\log p_{\bm{\theta}}(\bm{x}_{0}|\bm{z})$ is the Gaussian log-likelihood, which up to multiplicative constants, simplifies to:
$\log p_{\bm{\theta}}(\bm{x}_{0}|\bm{z})=\left\lVert\bm{x}_{0}-\bm{f}_{\bm{\theta}}(\bm{z})\right\rVert^{2}$ (45)
where $\bm{z}$ is a sample obtained from the posterior distribution with the reparametrization trick:
$\bm{z}=\bm{g}_{\bm{\phi}}^{\mu}(\bm{x}_{0})+\bm{g}_{\bm{\phi}}^{\sigma}(\bm{x}_{0})\bm{\epsilon}$ (46)
In discretized CT, for the triangular inequality, we have that:
$\left\lVert\bm{x}_{0}-\bm{f}_{\bm{\theta}}(\bm{z},1)\right\rVert^{2}\leq\sum_{i=0}^{N}\left\lVert\bm{f_{\theta}}(\bm{\psi}_{t_{i+1}}(\bm{x}_{0};\bm{x}_{1}),t_{i+1})-\bm{f}_{\bm{\theta}^{-}}(\bm{\psi}_{t_{i}}(\bm{x}_{0};\bm{x}_{1}),t_{i})\right\rVert^{2}.$ (47)
It is now easy to see that the consistency loss with variational noise coupling is also a lower bound on the evidence:
$\log p(\bm{x}_{0})$ $\leq\mathbb{E}_{q(\bm{z}|\bm{x}_{0})}\left[\left\lVert\bm{x}_{0}-\bm{f}_{\bm{\theta}}(\bm{z})\right\rVert^{2}\right]-\text{KL}(q(\bm{z}|\bm{x}_{0})\|p(\bm{z}))$ (48)
$\leq\mathbb{E}_{q(\bm{z}|\bm{x}_{0}),p(t_{i})}\left[\left\lVert\bm{f_{\theta}}(\bm{\psi}_{t_{i+1}}(\bm{x}_{0};\bm{x}_{1}),t_{i+1})-\bm{f}_{\bm{\theta}^{-}}(\bm{\psi}_{t_{i}}(\bm{x}_{0};\bm{x}_{1}),t_{i})\right\rVert^{2}\right]-\text{KL}(q(\bm{z}|\bm{x}_{0})\|p(\bm{z}))$ (49)
@@ C.1 Baselines
Here we recap the detials about the two baselines used in this work, the Improved Consistency Training from (Song & Dhariwal, 2024) and Easy Consistency Models from (Geng et al., 2024).
@@ C.2 Training details
Appendix D Toy experiments

### v2

@@ 7.1 Baselines and models
As baselines, we re-implement the iCT and ECM models, corresponding to our iCT-VE and ECM-VE. As an additional model, we add CT with the minibatch Optimal Transport Coupling (-OT) proposed in (Pooladian et al., 2023; Tong et al., 2024) and used in (Issenhuth et al., 2025), to compare the effectiveness and scalability of the OT coupling with the learned one. Finally, we combine the baselines with our proposed Variational Coupling (-VC). For the models with learned coupling, we use gradient clipping with a large value (200 in all the experiments) to avoid instabilities at the early stages of training.
@@ 7.2 Ablation for different $\beta$
To see the effect of $\beta$ on the generation performance, we compare the results for different values on the FashionMNIST dataset for iCT-VE-VC in Table 1. As expected, with small values of $\beta$, the coupling distribution deviates from the sampling distribution and the performance degenerates, while increasing $\beta$ to high values reduces the benefits of the learned coupling. More detailes on how we tune $\beta$ are reported in Appendix C.3.
@@
Table 1: Comparison of FID performance (lower is better) for one and two sampling steps, for varying values of $\beta$. The models are iCT-VE-VC and trained on the FashionMNIST dataset with the same settings described in appendix C. Best entries in bold.
@@
ECM-VE-VC (VCT) 3.26 / 2.02 5.47 / 4.16 5.08 / 3.15
ECM-LI-VC (VCT) 3.39 / 2.09 5.57 / 4.29 4.93 / 3.07
Table 3: Comparison of FID (lower is better, reported as 1-step / 2-step performance) for different models based on ECM. The model marked with a * is the baseline as reported in (Geng et al., 2025). All the other models are from our re-implementation. The best entries are highlighted in bold. For ImageNet, the results marked with $\dagger$ are obtained with models trained for $100$k iterations, while the others use $200$k iterations. Comprehensive comparisons with additional baselines are presented in Tables 9 and 10.
@@ 7.3 Results
CIFAR-10: For all the CIFAR-10 experiments, we use the DDPM++ architecture from (Song et al., 2021b) as implemented in (Karras et al., 2022), with EMA rate $0.9999$ as in (Geng et al., 2025). While this differs from the settings in (Song & Dhariwal, 2024), we found it to work better in our re-implementation. The remaining hyperparameters are the same as used in the respective baselines. From the results, we can see how using the learned coupling results in improved performance for both one and two steps generation, outperforming all the re-implemented baselines. The 1-step result from the original iCT is superior to our model. However, to the best of our knowledge, there is no open-source implementation that can reproduce the results reported in the paper. The learned coupling outperforms the minibatch OT coupling in all cases, as it is less affected by the effective (per device) batch size and the data dimensionality. Finally, our 2-step sampling performance for ECM-VE-VC is on par with the current SoTA achieved by other methods with similar settings (Wang et al., 2024; Lee et al., 2025; Lu & Song, 2025). We empirically compare the variance of the gradients for iCT-VE and iCT-VE-VC, and show in Figure 3 how the resulting reduced variance corresponds to improved FID score. In particular, in early training the model with VC exhibits higher variance, which can be attributed to the fact that the encoder is still learning the coupling. As training continues, the coupling becomes effective at providing better data-noise pairs to the model, which results in reduced gradient variance and improved generative performance.
@@
ImageNet $64\times 64$ (class conditional): As a baseline, we reuse the settings from ECM with the EDM2-S architecture and batch size $128$. While the baseline is trained for $100k$ iteration, we found that our models with Variational Coupling needed more time to converge properly, as the encoder weights are not pretrained and initialized at random. We therefore train our re-implemented baselines and models for $200k$ iterations instead, while we report the performance of our models trained for $100k$ iterations in Appendix C.5. In the $200k$ case, the models with Variational Coupling outperform the other models, with the LI kernel obtaining the best overall FID, while the OT coupling performs poorly due to the small batch size and high data dimensionality. In Figure 2 we compare samples from ECM-LI and ECM-LI-VC, where we can see how the images generated with VC are more clear and detailed.
@@ 8 Conclusions
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgment
We sincerely thank the reviewers for their constructive feedback, which greatly helped improve this work. We are also deeply grateful to our colleague, Naoki Murata, for revising the manuscript and providing valuable comments.
@@ Appendix B Consistency Lower Bound
Proposition B.1.
The following upper bound for negative log-density holds:
$-\log p_{\theta}(\bm{x}_{0})$ $\leq\frac{1}{2\sigma^{2}}\,\mathbb{E}_{q_{\phi}(\bm{x}_{1}\mid\bm{x}_{0})}\|\bm{x}_{0}-\bm{f_{\theta}}(\bm{x}_{1},1)\|^{2}+\mathcal{D}_{\mathrm{KL}}\Bigl{(}q_{\phi}(\bm{z}\mid\bm{x}_{0})\,\Big{\|}\,p(\bm{z})\Bigr{)}$
@@
This proposition establishes the connection between the minimization objective of CT-VC and that of the continuous-time CM.
Proof.
For the triangle inequality, we have:
$\left\lVert\bm{x}_{0}-\bm{f}_{\bm{\theta}}(\bm{x}_{1},1)\right\rVert\leq$ (41)
$\sum_{i=0}^{N}\left\lVert\bm{f_{\theta}}(\bm{\psi}_{t_{i+1}}(\bm{x}_{0};\bm{x}_{1}),t_{i+1})-\bm{f}_{\bm{\theta}^{-}}(\bm{\psi}_{t_{i}}(\bm{x}_{0};\bm{x}_{1}),t_{i})\right\rVert.$
We can now square both sides, obtaining:
$\left(\sum_{i=0}^{N}\left\lVert\bm{f_{\theta}}(\bm{\psi}_{t_{i+1}}(\bm{x}_{0};\bm{x}_{1}),t_{i+1})-\bm{f}_{\bm{\theta}^{-}}(\bm{\psi}_{t_{i}}(\bm{x}_{0};\bm{x}_{1}),t_{i})\right\rVert\right)^{2}.$
Now, for the Cauchy-Schwarz Inequality, we can write the right hand side as:
$\left(\sum_{i=0}^{N}\left\lVert\bm{f_{\theta}}(\bm{\psi}_{t_{i+1}}(\bm{x}_{0};\bm{x}_{1}),t_{i+1})-\bm{f}_{\bm{\theta}^{-}}(\bm{\psi}_{t_{i}}(\bm{x}_{0};\bm{x}_{1}),t_{i})\right\rVert\right)^{2}\leq$ (43)
$N\sum_{i=0}^{N}\left\lVert\bm{f_{\theta}}(\bm{\psi}_{t_{i+1}}(\bm{x}_{0};\bm{x}_{1}),t_{i+1})-\bm{f}_{\bm{\theta}^{-}}(\bm{\psi}_{t_{i}}(\bm{x}_{0};\bm{x}_{1}),t_{i})\right\rVert^{2}.$
Since:
$\log p_{\theta}(\bm{x}_{0}\mid\bm{x}_{1})\propto-\frac{1}{2\sigma^{2}}\|\bm{x}_{0}-\bm{f_{\theta}}(\bm{x}_{1},1)\|^{2},$ (44)
we can write a lower bound on the log density as:
$\log p_{\theta}(\bm{x}_{0})\geq-\frac{1}{2\sigma^{2}}\,\mathbb{E}_{q_{\phi}(\bm{x}_{1}\mid\bm{x}_{0})}\|\bm{x}_{0}-\bm{f_{\theta}}(\bm{x}_{1},1)\|^{2}-\mathcal{D}_{\mathrm{KL}}\Bigl{(}q_{\phi}(\bm{x}_{1}\mid\bm{x}_{0})\,\Big{\|}\,p(\bm{x}_{1})\Bigr{)}.$ (45)
In summary, the ELBO bound for the Gaussian VAE is given by
$-\log p_{\theta}(\bm{x}_{0})\leq\frac{1}{2\sigma^{2}}\,\mathbb{E}_{q_{\phi}(\bm{x}_{1}\mid\bm{x}_{0})}\|\bm{x}_{0}-\bm{f_{\theta}}(\bm{x}_{1},1)\|^{2}+\mathcal{D}_{\mathrm{KL}}\Bigl{(}q_{\phi}(\bm{x}_{1}\mid\bm{x}_{0})\,\Big{\|}\,p(\bm{x}_{1})\Bigr{)}:=L(\bm{x}_{0};\theta,\phi).$ (46)
Combining the inequalities we derived in Eq. 42 and 43, below we consider the case when $N\rightarrow\infty$. First, we define:
$S_{N}=N\sum_{i=0}^{N}\mathbb{E}_{q(x_{1}\mid x_{0}),p(t_{i})}\left[\|\bm{f_{\theta}}(\psi_{t_{i+1}}(x_{0};x_{1}),t_{i+1})-\bm{f_{\theta}}(\psi_{t_{i}}(x_{0};x_{1}),t_{i})\|^{2}\right].$ (47)
We assume that $t_{i}$ is a partition of the interval $[0,1]$ of:
$\Delta t:=t_{i+1}-t_{i}=\mathcal{O}\left(\frac{1}{N}\right).$ (48)
For small $\Delta t=\mathcal{O}\left(\frac{1}{N}\right)$, we approximate the squared difference in function values using a first-order Taylor expansion [^1] Here, we assume $\psi_{t_{i+1}}-\psi_{t_{i}}=\mathcal{O}(t_{i+1}-t_{i})$:
$\|\bm{f_{\theta}}(\psi_{t_{i+1}},t_{i+1})-\bm{f_{\theta}}(\psi_{t_{i}},t_{i})\|^{2}\approx\left\|\frac{d}{dt}\bm{f_{\theta}}(\psi_{t},t)\bigg{|}_{t=t_{i}}\right\|^{2}\cdot\Delta t^{2}.$ (49)
Since $\Delta t=\mathcal{O}\left(\frac{1}{N}\right)$, we substitute:
$\|\bm{f_{\theta}}(\psi_{t_{i+1}},t_{i+1})-\bm{f_{\theta}}(\psi_{t_{i}},t_{i})\|^{2}\approx\frac{1}{N^{2}}\left\|\frac{d}{dt}\bm{f_{\theta}}(\psi_{t},t)\right\|^{2}.$ (50)
Multiplying by $N$:
$S_{N}=N\sum_{i=0}^{N}\mathbb{E}\left[\|\bm{f_{\theta}}(\psi_{t_{i+1}},t_{i+1})-\bm{f_{\theta}}(\psi_{t_{i}},t_{i})\|^{2}\right]\approx\frac{1}{N}\sum_{i=0}^{N}\mathbb{E}\left[\left\|\frac{d}{dt}\bm{f_{\theta}}(\psi_{t},t)\right\|^{2}\right].$ (51)
As $N\rightarrow\infty$,
@@
Thus, multiplying by $N$, we obtain:
[…]
Combining the above limit with the ELBO bound:
$-\log p_{\theta}(\bm{x}_{0})$ $\leq\frac{1}{2\sigma^{2}}\,\mathbb{E}_{q_{\phi}(\bm{x}_{1}\mid\bm{x}_{0})}\|\bm{x}_{0}-\bm{f_{\theta}}(\bm{x}_{1},1)\|^{2}+\mathcal{D}_{\mathrm{KL}}\Bigl{(}q_{\phi}(\bm{z}\mid\bm{x}_{0})\,\Big{\|}\,p(\bm{z})\Bigr{)}$ (54)
$\leq\frac{1}{2\sigma^{2}}\frac{1}{N}\sum_{i=0}^{N}\mathbb{E}\left[\left\|\frac{d}{dt}\bm{f_{\theta}}(\psi_{t},t)\right\|^{2}\right]+\mathcal{D}_{\mathrm{KL}}\Bigl{(}q_{\phi}(\bm{z}\mid\bm{x}_{0})\,\Big{\|}\,p(\bm{z})\Bigr{)}.$ (55)
[…]
@@ C.1 Baselines
Here we recap the detials about the two baselines used in this work, the Improved Consistency Training from (Song & Dhariwal, 2024) and Easy Consistency Models from (Geng et al., 2025).
@@ C.2 Training details
C.3 Tuning $\beta$
In our experiments, we tuned $\beta$ with a coarse grid search with different values with a gap of 10. For iCT on CIFAR10, we initially tested the values $\beta=[10,20,30,40]$, of which $\beta=30$ gave the best performance, then tested also for $\beta=[25,35]$ which did not improve the performance. Similarly, for ECM on CIFAR10, we tested for $\beta=[10,20,30,40]$, and after achieving the best performance with $\beta=10$, we tested $\beta=[5,15]$ which did not improve the performance. The tuning was done with the VE kernel and the best values were used also for the LI kernel. We used the best values of $\beta$ also in FashinMNIST and FFHQ without additional tuning. For ImageNet we observed on early runs that $\beta$ needed to be much larger, so we initially tuned for $\beta=[30,60,90,120]$. After achieving the best results with the VE kernel for $\beta=90$, we further tuned for $\beta=[70,80,90,100,110]$ for both VE and LI kernels, and found $\beta=100$ to be the best for VE and $\beta=90$ for LI.
C.4 The effect of gradient clipping
As our method uses gradient clipping for stability, we test the effect of gradient clipping on the baselines for a fair comparison. We report the results in tables 7 and 7 for iCT and ECM respectively. While the results generally improve slightly in iCT based methods, it is cleare that the main performance gain in our method is given by the improved coupling.
Method 1-step 2-step
iCT-VE (w/o gc) 3.61 2.79
iCT-VE (w/ gc) 3.49 2.57
iCT-LI (w/o gc) 3.81 2.87
iCT-LI (w/ gc) 3.54 2.66
iCT-VE-OT (w/o gc) 3.28 2.66
iCT-VE-OT (w/ gc) 3.21 2.56
iCT-LI-OT (w/o gc) 3.42 2.77
iCT-LI-OT (w/ gc) 3.18 2.63
iCT-VE-VC (w/ gc) (VCT) 2.86 2.32
iCT-VE-LI (w/ gc) (VCT) 2.94 2.32
Table 6: FID performance with/without gc on CIFAR10 for iCT based methods.
Method 1-step 2-step
ECM-VE (w/o gc) 3.68 2.14
iCT-VE (w/ gc) 3.70 2.12
iCT-LI (w/o gc) 3.65 2.14
iCT-LI (w/ gc) 3.76 2.17
iCT-VE-OT (w/o gc) 3.46 2.13
iCT-VE-OT (w/ gc) 3.45 2.12
iCT-LI-OT (w/o gc) 3.49 2.13
iCT-LI-OT (w/ gc) 3.50 2.12
iCT-VE-VC (w/ gc) (VCT) 3.26 2.02
iCT-VE-LI (w/ gc) (VCT) 2.39 2.09
Table 7: FID performance with/without gc on CIFAR10 for ECM based methods.
C.5 Imagenet training with $100k$ iterations
We report here the ImageNet experiments with 100k iterations. For ECM-LI-VC we increased $\beta$ to $\beta=100$ as $\beta=90$ diverged during training. Note that for runs with $100k$ iterations, we sometimes encountered divergences for our models with small $\beta$ and sometimes also for the baselines, while it seems to be solved when training for 200k iterations. The 1-step/2-step FID results are reported in Table 8. For the settings with 100k iterations, our method performs similarly or slightly worse than the baseline. We believe this is due to the fact that the encoder for our model requires more iterations to learn the coupling, as demonstrated by the improved results for 200k iterations.
Method 1-Step FID 2-Step FID
ECM-VE 5.66 3.78
ECM-LI 5.63 3.48
ECM-VE-OT 6.74 4.71
ECM-LI-OT 6.65 4.64
ECM-VE-VC (VCT) 5.67 3.67
ECM-LI-VC (VCT) 6.34 3.77
Table 8: Results of our models trained on class conditional IageNet $64\times 64$ for $100k$ iterations with batch size 128.
Appendix D Comprehensive comparison
In this section we report a comprehensive comparison with other methods on CIFAR10 and ImageNet $64\times 64$, in tables 9 and 10.
Table 9: FID, NFE and # param. on CIFAR-10. Bold indicates the best result for each category and NFE.
Method NFE FID # param. (M)
Diffusion models
EDM (Karras et al., 2022) 35 1.97 55.7
PFGM++ (Xu et al., 2023) 35 1.91 55.7
DDPM (Ho et al., 2020) 1000 3.17 35.7
LSGM (Vahdat et al., 2021) 147 2.10 475
Consistency models
   1‑step
iCT (Song & Dhariwal, 2024) 1 2.83 56.4
iCT‑deep (Song & Dhariwal, 2024) 1 2.51 112
CTM (Kim et al., 2024) (w/o GAN) 1 5.19 55.7
ECM (Geng et al., 2025) 1 3.60 55.7
TCM (Lee et al., 2025) 1 2.46 55.7
sCT (Lu & Song, 2025) 1 2.85 -
iCT-VE-VC (VCT) 1 2.86 57.3
iCT-LI-VC (VCT) 1 2.94 57.3
ECM-VE-VC (VCT) 1 3.26 57.3
ECM-LI-VC (VCT) 1 3.39 57.3
   2‑step
iCT (Song & Dhariwal, 2024) 2 2.46 56.4
iCT‑deep (Song & Dhariwal, 2024) 2 2.24 112
ECM (Geng et al., 2025) 2 2.11 55.7
TCM (Lee et al., 2025) 2 2.05 55.7
sCT (Lu & Song, 2025) 2 2.06 -
iCT-VE-VC (VCT) 2 2.32 57.3
iCT-LI-VC (VCT) 2 2.32 57.3
ECM-VE-VC (VCT) 2 2.02 57.3
ECM-LI-VC (VCT) 2 2.09 57.3
Variational score distillation
DMD (Yin et al., 2024) 1 3.77 55.7
Diff‑Instruct (Luo et al., 2023) 1 4.53 55.7
SiD (Zhou et al., 2024) 1 1.92 55.7
Knowledge distillation
KD (Luhman & Luhman, 2021) 1 9.36 35.7
DSNO (Zheng et al., 2023) 1 3.78 65.8
TRACT (Berthelot et al., 2023) 1 3.78 55.7
2 3.32 55.7
PD (Salimans & Ho, 2022) 1 9.12 60.0
2 4.51 60.0
CD (LPIPS) (Song et al., 2023) 1 3.55 -
2 2.93 -
sCD (Lu & Song, 2025) 1 3.66 -
2 2.52 -
Table 10: FID, NFE and ##\## param. on ImageNet $64\times 64$. Dotted lines separate results by ##\## param. Bold indicates the best result for each category and NFE. Our models, marked with a , use batch size 128 while equivalent models based on ECM use batch size 1024.
Method NFE FID # param. (M)
Diffusion models
   EDM2‑S (Karras et al., 2024) 63 1.58 280
   EDM2‑XL (Karras et al., 2024) 63 1.33 1119
Consistency models
   1‑step
   iCT (Song & Dhariwal, 2024) 1 4.02 296
   iCT‑deep (Song & Dhariwal, 2024) 1 3.25 592
   ECM (Geng et al., 2025) (EDM2‑S) 1 4.05 280
   TCM (Lee et al., 2025)(EDM2‑S) 1 2.88 280
   sCT (Lu & Song, 2025) (EDM2-S) 1 3.23 280
   ECM (Geng et al., 2025) (EDM2‑XL) 1 2.49 1119
   TCM (Lee et al., 2025) (EDM2‑XL) 1 2.20 1119
   sCT (Lu & Song, 2025) (EDM2-XL) 1 2.04 1119
   ECM-VE-VC (Ours, EDM2-S) 1 5.08 286
   ECM-LI-VC (Ours, EDM2-S) 1 4.93 286
   2‑step
   iCT (Song & Dhariwal, 2024) 2 3.20 296
   iCT‑deep (Song & Dhariwal, 2024) 2 2.77 592
   ECM (Geng et al., 2025) (EDM2‑S) 2 2.79 280
   TCM (Lee et al., 2025)(EDM2‑S) 2 2.31 280
   sCT (Lu & Song, 2025) (EDM2-S) 2 2.93 280
   ECM (Geng et al., 2025)(EDM2‑XL) 2 1.67 1119
   TCM (Lee et al., 2025)(EDM2‑XL) 2 1.62 1119
   sCT (Lu & Song, 2025) (EDM2-XL) 2 1.48 1119
   ECM-VE-VC (Ours, EDM2-S) 2 3.15 286
   ECM-LI-VC (Ours, EDM2-S) 2 3.07 286
Variational score distillation
   DMD2 w/o GAN (Yin et al., 2024) 1 2.60 296
   Diff‑Instruct (Luo et al., 2023) 1 5.57 296
   EMD‑16 (Xie et al., 2024) 1 2.20 296
   Moment Matching (Salimans et al., 2024) 1 3.00 400
2 3.86 400
   SiD (Zhou et al., 2024) 1 1.52 296
Knowledge distillation
   DSNO (Zheng et al., 2023) 1 7.83 329
   TRACT (Berthelot et al., 2023) 1 7.43 296
2 4.97 296
   PD (Salimans & Ho, 2022) 1 15.4 296
2 8.95 296
   CD (LPIPS) (Song et al., 2023) 1 6.20 -
2 4.70 -
   sCD (Lu & Song, 2025) (EDM2-XL) 1 2.44 1119
2 1.66 1119
Appendix E Toy experiments